# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

### v2

@@ Appendix B Other experiments
Wikitext Word Perplexity $\downarrow$ 204.94 141.56 357.95 213.74
Table 7: Passkey retrieval results including the other experiments.
Model Sequence Length
890 1983 3075 4167 4986
Softmax 95.5 97.0 73.0 70.5 0.0
Softpick 94.0 91.0 75.0 66.5 0.0
Rectified-Only 95.5 68.0 57.5 24.5 0.0
Scalable-Softmax 98.0 96.9 87.0 55.0 0.0
Appendix C Full Result on Quantization
Table 8: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@ Appendix C Full Result on Quantization
Table 9: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Table 10: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Appendix D Full Result on Quantization on 1.8B
Table 11: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Table 12: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = Softpick - Softmax.
@@
Table 13: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Appendix E More Attention Maps

### v3

@@ Appendix E Passkey Retrieval Results
Wikitext Word Perplexity $\downarrow$ 24.01 24.41 27.93 25.13
Appendix F Full Results on Quantization of the 340M Models
Table 7: Comparison of softpick vs softmax performance for BNB Dettmers et al. (2022, 2023) quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@ Appendix F Full Results on Quantization of the 340M Models
Table 8: Comparison of softpick vs softmax performance for GPTQ Frantar et al. (2022) quantization. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Table 9: Comparison of softpick vs softmax performance for HQQ Badri and Shaji (2023) quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Appendix G Full Results on Quantization of the 1.8B Models
Table 10: Comparison of softpick vs softmax performance for BNB Dettmers et al. (2022, 2023) quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Table 11: Comparison of softpick vs softmax performance for GPTQ Frantar et al. (2022) quantization. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = Softpick - Softmax.
@@
Table 12: Comparison of softpick vs softmax performance for HQQ Badri and Shaji (2023) quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
@@
Appendix H More Attention Maps